Please help me with this problem:
What is the equation for the asymptote for:
R(x)=x^3+1/x^2+1

Found 2 solutions by jim_thompson5910, stanbon:
Answer by jim_thompson5910(21685)

Perform polynomial long division like this:

Photobucket


Since the quotient is x, this means that the equation of the asymptote is y=x

Answer by stanbon(48558)

Please help me with this problem:
What is the equation for the asymptote for:
R(x)=x^3+1/x^2+1
------------
Vertical asymptote: None because the denominator is never zero.
--------------
Horizontal asymptote: 1x^3/0x^3 is undefined; so no horizontal asymptote
---------------
Slant asymptote:
R(x)/(x^2+1) = (x-1) + 2/(x^2+1)
Slant asymptote is y = x-1
==============================
Cheers,
Stan H.